Database last updated on Tue, 03 Feb 2026 06:10:53 +0000 by localhost

Critical warnings from b-files' comments for days between 2023167 and 2023167 (0 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
hp:Spectrometer alignment trouble.11B16723.194Toron#19400:50:061-1
rl:2023 165 01 11 13 Motor #9: appears stuck with a reference sensor activated.31B16723.194Toron#19405:26:051-1
rl:2023 165 05 23 25 Motor 11: Move attempted outside of logical bounds11B16723.194Toron#19405:26:381-1
dt:Brewer failed to respond 5 times21B16723.194Toron#19408:27:301-1
hg:Lamp not on... hg test terminated F1= 1 F4= 011B16723.194Toron#19413:09:171-1
hg:Lamp not detected f1= 1 f4= 011B16723.194Toron#19413:09:181-1

Instrument constant changes for the week before 2023167 (3 sec, 0 rows)

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
19420231671146.9

Non-critical warnings from b-files' comments for days between 2023167 and 2023167 (0 sec, 11 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
ds:DS intensity too low ( 2/cy), skipping.299B16723.194Toron#19401:17:111-1
zs:ZS intensity too low ( 31/cy), skipping.299B16723.194Toron#19401:17:261-1
rl:Log item overwritten by later entry.23B16723.194Toron#19405:26:031-1
rl:2023 165 01 11 13 Motor #9: appears stuck with a reference sensor activated.13B16723.194Toron#19405:26:051-1
rl:2023 165 01 11 13 Motor #10: appears stuck with a reference sensor activated.23B16723.194Toron#19405:26:071-1
rl:2023 165 02 15 55 Bad command string. Command string buffer cleared.33B16723.194Toron#19405:26:261-1
rl:2023 165 02 15 55 Bad command string. Command string buffer cleared.33B16723.194Toron#19405:26:261-1
rl:2023 165 02 15 55 <33B16723.194Toron#19405:26:281-1
rl:2023 165 04 51 58 New value set for watermark SUPPLY.VOLTAGE.HIGH[23B16723.194Toron#19405:26:301-1
sc:SC intensity too low ( 293/cy), skipping.43B16723.194Toron#19412:40:501-1
ds:Suppressed oscillation of filter wheel.45B16723.194Toron#19415:44:391-1

Measurements and tests summary.


_jday_loc_serial_skcaodedshghpsluvzs
167Toronto#194calsc,Must do HG now,Next schedule will be calsc (now calsc ZS= 90 ZS= 111.806),Next schedule will be calsc (now calsc ZS= 90 ZS= 111.949),Next schedule will be calsc (now calsc ZS= 90 ZS= 91.189),Next schedule will be calsc (now calsc ZS= 90 ZS= 91.753)657334448139
(0 sec)

Relative and absolute humidity for day 2023167 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
1947.82.2